OpenStudy (anonymous):

Can you help me find the arc length of a pizza with 10 slices and diameter of 12”? Here’s the interior angle: Interior angle = 360°/10 = 36° And this is what I've got so far, what do i do next.. Arc length = 36°/360°(π•10) Arc length = 0.1 (π • 10)

OpenStudy (anonymous):

is this just you search the circular of pizza the divide into 10?

OpenStudy (anonymous):

huh..?

hartnn (hartnn):

diameter, d= 12 , so it would be Arc length = 36°/360°(π•12) isn't it ?

OpenStudy (anonymous):

@hartnn i was actually just about you ask you to help me, and now i see you're viewing this haha Arc length = 0.1 (π • 10) = ___ for the other question it was Arc length = 0.0833333333333333 (π • 12) = 3.1414 inch but i don't know how you got that

OpenStudy (anonymous):

and i thought it would be (π•10) because it's 10 slices now

hartnn (hartnn):

Arc length = a/360°(πd) d is the diameter = 12

OpenStudy (amistre64):

circumference divided by slices :)

OpenStudy (amistre64):

as long as all the slices are the same size

OpenStudy (amistre64):

i see the OS is shuffling the posts :/

hartnn (hartnn):

even i do .... :/

OpenStudy (anonymous):

wait so did i do something wrong? eek

hartnn (hartnn):

u know this formula , right ? Arc length = a/360°(πd)

hartnn (hartnn):

u put a= 36, d=12 and u ar e done

hartnn (hartnn):

Arc length = 0.1 (π • 12)

OpenStudy (anonymous):

But when I said my pizza had 12 slices and was 12” in diameter the arc length was 0.0833333333333333 (π • 12) = 3.1414 inch So now it’s 10 slices and still the whole thing is 12” in diameter so the arc length is Arc length = 0.1 (π • 10) = ________ i'm just so confused

hartnn (hartnn):

d is diameter, not the number of slices

OpenStudy (anonymous):

OH okay hahaha so what do i write instead then?

hartnn (hartnn):

Arc length = 36°/360°(π•12) Arc length = 0.1 (π • 12)= 3.7399 inch

OpenStudy (anonymous):

you are a god omg

OpenStudy (anonymous):

thank you SO MUCH YOU HAVE HELPED ME OUT LIKE CRAZY

hartnn (hartnn):

i assume u meant 'good' welcome ^_^

OpenStudy (anonymous):

haha